# Visualising Policy-Reward Interplay to Inform Zeroth-Order Preference Optimisation of Large Language Models

### v1

@@ Abstract
Fine-tuning LLMs with first-order methods like back-propagation is computationally intensive. Zeroth-Order (ZO) optimisation, using function evaluations instead of gradients, reduces memory usage but suffers from slow convergence in high-dimensional models. As a result, ZO research in LLMs has mostly focused on classification, overlooking more complex generative tasks. In this paper, we introduce ZOPrO, a novel ZO algorithm designed for Preference Optimisation in LLMs. We begin by analysing the interplay between policy and reward models during traditional (first-order) Preference Optimisation, uncovering patterns in their relative updates. Guided by these insights, we adapt Simultaneous Perturbation Stochastic Approximation (SPSA) with a targeted sampling strategy to accelerate convergence. Through experiments on summarisation, machine translation, and conversational assistants, we demonstrate that our method consistently enhances reward signals while achieving convergence times comparable to first-order methods. While it falls short of some state-of-the-art methods, our work is the first to apply Zeroth-Order methods to Preference Optimisation in LLMs, going beyond classification tasks and paving the way for a largely unexplored research direction. Code and visualisations are available at https://github.com/alessioGalatolo/VisZOPrO.
@@ 1 Introduction
This has spurred interest in alternative optimisation techniques, such as Zeroth-Order Optimization Malladi et al. (2023); Zhang et al. (2024); Shirkavand et al. (2025). Zeroth-Order optimisation (Figure 1, left) presents a compelling alternative by eliminating the need for gradient computation. Instead, it approximates gradients using only function evaluations (i.e., forward passes), typically through finite difference methods. For example, in Simultaneous Perturbation Stochastic Approximation (SPSA) Spall (1992); Malladi et al. (2023) the function is evaluated two times, where each time noise is either added or subtracted from the weights and a ‘step’ is taken toward the most promising direction. This significantly reduces memory requirements, as the model needs to be loaded into memory only once. Additionally, due to their very nature, these methods are also inherently fit for optimising non-differentiable objectives such as human preferences Malladi et al. (2023) where a step can be taken directly towards the preferred model’s generations. While promising, Zeroth-Order methods face challenges in terms of convergence speed, particularly for high-dimensional models like LLMs.
@@
In contrast to previous work, we dive into the more challenging and practically relevant problem of Preference Optimisation for LLMs. Preference optimisation involves training a policy model to generate desirable outputs, often guided by a reward model that learns to distinguish between high-quality and low-quality outputs. This process in many cases entails iterative training of both models, generally in settings of online learning Guo et al. (2024a) or as a way to refine the reward model to account for the distribution shift of the policy throughout training Wang et al. (2024).
To this end, we hypothesise that the evolving dynamics of the policy and reward models during Preference Optimisation can provide valuable information for guiding the search direction of Zeroth-Order methods. We thus begin our investigation by visualising and analysing how popular preference optimisation methods, such as Proximal Policy Optimization (PPO) Schulman et al. (2017); Ouyang et al. (2022) (and variations, Ahmadian et al., 2024) and Direct Preference Optimization [^1] As standard DPO does not use any reward directly, we choose one of its variations Guo et al. (2024a) where we can integrate the reward model. (DPO) Rafailov et al. (2023), act on the policy and reward model, with particular attention on their relation. Whilst we do not find any statistical direct correlation, our visualisation and analysis suggest a significant influence of previous updates on upcoming ones. Further, we find that reward model’s updates are consistently orthogonal to those of the policy.
@@
    Based on the observed dynamics, we develop ZOPrO, a Zeroth-Order Optimization algorithm specifically tailored for preference optimisation in LLMs. ZOPrO incorporates a novel mechanism to guide the search direction by utilising the inter-dependencies between the policy and reward models.
@@
    We conduct experiments on multiple tasks-datasets to demonstrate the effectiveness of ZOPrO in fine-tuning LLMs for Preference Optimisation. Our results show that ZOPrO achieves comparable performance to first-order methods while significantly reducing memory usage and iteration time.
@@ 2.1 Preference Optimisation
Reinforcement Learning from Human Feedback (RLHF), or more generally Preference Optimisation, aims to optimise Machine Learning models to align with complex human judgements and preferences Christiano et al. (2017). Typically, human feedback data is used to train a reward model, often conceptualised with the Bradley-Terry model Bradley and Terry (1952), as an effective approach to converting human preferences into continuous reward signals (Bai et al., 2022). The reward model is then used to optimise a policy model using Reinforcement Learning algorithms, such as Proximal Policy Optimization (PPO) Schulman et al. (2017). RLHF has been applied in LLMs across various tasks from sentiment generation Ziegler et al. (2019) to summarisation Stiennon et al. (2020b) to instruction following Ouyang et al. (2022). In this framework, the policy and reward models influence each other in a shared interplay. Updates to the policy shift the text distribution that the reward model evaluates (Lu et al., 2024; Wang et al., 2024; Dou et al., 2024), while updates to the reward model alter the policy’s optimisation trajectory (Li et al., 2024).
Newer methods, such as Direct Preference Optimization (DPO) Rafailov et al. (2023) and variations Liu et al. (2024); Gheshlaghi Azar et al. (2024); Ji et al. (2024) avoid the use of Reinforcement Learning, instead relying on indirect reward. However, other works suggest the use of a reward model or external judge can still be beneficial for the training Guo et al. (2024a). Whilst promising, these methods have not replaced RL-based approaches. Newer RL approaches include Group Relative Policy Optimization (GRPO) Shao et al. (2024) which introduces group-based relative rewards to accelerate convergence and obtain substantial improvements in math and general reasoning DeepSeek-AI (2025) or Ahmadian et al. (2024) which introduces REINFORCE Leave-One-Out (RLOO), a REINFORCE-based approach competitive to PPO with gains in memory utilisation and convergence speed.
@@ 3 Visualising Preference Optimisation
We begin our investigation by analysing how the policy and reward models are optimised using traditional methods. We simulate multiple training runs across different methods, collecting model checkpoints at various stages of optimisation. At the end of training, we visualise and analyse each method with particular attention to the interplay between the reward and policy models.
@@ 3.1 Experimental Setup
We select the two most used methods for preference optimisation: PPO Schulman et al. (2017) and DPO Rafailov et al. (2023); for the latter we select its online version Guo et al. (2024a) which integrates a reward model—an essential component for our analysis. The models are trained using Anthropic’s HH-RLHF dataset Bai et al. (2022), chosen due to its widespread adoption in preference modelling tasks. After initial testing, we report convergence difficulties with PPO, where the model easily collapses into generating only eos tokens. Due to this, we analyse one of its variations, RLOO Ahmadian et al. (2024) instead.
For all methods, we adopt an iterative training regime in which, after each iteration, the reward model is updated to better evaluate samples from the policy model, accounting for distribution shifts. Our algorithm is shown in 1.
@@ 3.3 Results
We show some 2D visualisation of different methods in Figure 3 and 3D visualisations in 2. We include interactive visualisations at https://alessiogalatolo.github.io/VisZOPrO/.
@@
These insights align with existing literature, where DPO has been shown to exhibit more stable training dynamics compared to RL-based approaches.
@@ 3.4 Quantitative Analysis
Motivated by the insights on the relative angles we do a more rigorous analysis on orthogonality. For this we use the weights in their original dimension and besides comparing the weights of the whole model(s), we also check layer-wise orthogonality. Our results suggest consistent orthogonality between the reward and policy updates across both methods. This aligns with prior findings that orthogonality stems from their distinct optimisation paradigms (Yu et al., 2020): the reward model, guided by supervised learning, predicts human preference scores, while the policy model, trained via Preference Optimisation, optimises text generation capabilities. Their gradients emerge from fundamentally different objectives—–discriminative in reward modelling and generative in policy optimisation—–driving orthogonal trajectories despite reward-model-guided updates.
@@ 4.1 Formalisation
Our proposed ZOPrO method utilises this structure in conjunction with SPSA Spall (1992), as the base of the ZO algorithm, and RLOO Ahmadian et al. (2024), as the objective function.
@@ 4.1.1 Optimisation via SPSA
Note, how our approach is also validated by previous attempts at surrogate or prior-guided strategies in ZO methods such as Cheng et al. (2021); Meier et al. (2019), where these previous methods assume to have multiple ‘oracle gradient directions’. In our case, we ground our perturbation strategy in empirical evidence and re-use the updates from the previous iteration. In particular, prior works had to use ‘oracle’ directions as ZO updates are very noisy locally and cannot be used directly. Our method instead leverages the weight updates over a whole iteration, making their use possible.
@@ 4.1.2 Adaptive Perturbation Strategy
To compute perturbation directions $z$, we first utilise $\Delta\pi_{t}$ with a scaling factor $\alpha$, which starts at 1 and decays to 0 within each iteration. Next, we sample a random noise vector from a multivariate standard normal distribution, $\mathcal{N}(0,I_{d})$, and project it onto the orthogonal complement of $\Delta R_{t}$, yielding $u_{i}$. A preliminary perturbation vector $\hat{z}_{i}$ is given by:
@@ 4.2 Convergence Analysis
First, the deterministic component $\Delta\pi_{t}$ in our perturbation formulation vanishes as $\alpha$ decays to zero, ensuring that stochastic exploration remains dominant in later iterations. This preserves the key stochastic approximation properties required for convergence.
@@ 5 Experiments
Among the challenges of effective optimisation of ZO methods is a careful balance of the hyper-parameters $\eta$ and $\varepsilon$: when these values are too high, the model rapidly degenerates into generating incoherent outputs; if they are too low, the model fails to improve. The challenge is further exacerbated in the context of Preference Optimisation, where improvements are subtle compared to classification tasks (previous works), in which performance gains are more easily quantifiable. For this, we sweep over multiple values of $\varepsilon$ (the perturbation multiplier) and learning rates $\eta$ using a standard grid search approach where we check 9 combinations of these values. We select the most promising ones based on their reward improvement and keep them fixed across tasks.
For our experiments, we adopt a similar iterative training to that described in Section 3.1. We consider a single model, LLama 3.2 1B, the previously used task of building a conversational assistant and add two new tasks: summarisation, through the dataset ‘Summarize from Feedback’ Stiennon et al. (2020a), and machine translation through the dataset WMT20 Barrault et al. (2020).
@@ 5.1 Results
We show in Table 1 some example generations after two iterations of training with our method. We can see that the policy appears proficient in all of the tasks that we considered. In Figure 4 we show various metrics comparing a training iteration of our method with RLOO and PPO [^3] We do not compare with Online DPO due to the difference in objectives between the methods.. Here, we report the average reward at the end of the iteration compared to the beginning, the time the method took to complete the iteration, the reward ‘gained’ over time (i.e., how fast in wall-clock time the method is improving) and the peak memory usage. We can see that our method whilst reporting the lowest reward ‘gains’ at the end of the iteration, is the fastest. This conforms to previous ZO methods where every step brings very little improvement but at much less cost in terms of time and memory.
@@
During our experiments, we also report particular difficulty in choosing the best values of $\varepsilon$ and $\eta$. This generally concretises in suboptimal training times and sometimes in policy degeneration. In our cases, we found it often beneficial to have a bigger learning rate $\eta$ than perturbation parameter $\varepsilon$, landing on $41e-$ 4 and $51e-$ 5 as optimal values.

### v2

@@ Abstract
Fine-tuning Large Language Models (LLMs) with first-order methods like back-propagation is computationally intensive. Zeroth-Order (ZO) optimisation uses function evaluations instead of gradients, reducing memory usage, but suffers from slow convergence in high-dimensional models. As a result, ZO research in LLMs has mostly focused on classification, overlooking more complex generative tasks. In this paper, we introduce ZOPrO, a novel ZO algorithm designed for Preference Optimisation in LLMs. We begin by analysing the interplay between policy and reward models during traditional (first-order) Preference Optimisation, uncovering patterns in their relative updates. Guided by these insights, we adapt Simultaneous Perturbation Stochastic Approximation (SPSA) with a targeted sampling strategy to accelerate convergence. Through experiments on summarisation, machine translation, and conversational assistants, we demonstrate that our method consistently enhances reward signals while achieving convergence times comparable to first-order methods. While it falls short of some state-of-the-art methods, our work is the first to apply Zeroth-Order methods to Preference Optimisation in LLMs, going beyond classification tasks and paving the way for a largely unexplored research direction. Code and visualisations are available at https://github.com/alessioGalatolo/VisZOPrO.
@@ 1 Introduction
This has spurred interest in alternative optimisation techniques, such as Zeroth-Order Optimisation Malladi et al. (2023); Zhang et al. (2024); Shirkavand et al. (2025). Zeroth-Order optimisation (Figure 1, left) presents a compelling alternative by eliminating the need for gradient computation. Instead, it approximates gradients using only function evaluations (i.e., forward passes), typically through finite difference methods. For example, in Simultaneous Perturbation Stochastic Approximation (SPSA) Spall (1992); Malladi et al. (2023), the function is evaluated two times, where each time noise is either added or subtracted from the weights and a ‘step’ is taken toward the most promising direction. This significantly reduces memory requirements, as the model needs to be loaded into memory only once. Additionally, due to their very nature, these methods are also inherently fit for optimising non-differentiable objectives such as human preferences Malladi et al. (2023) where a step can be taken directly towards the preferred model’s generations. While promising, Zeroth-Order methods face challenges in terms of convergence speed, particularly for high-dimensional models like LLMs.
@@
In contrast to previous work, we dive into the more challenging and practically relevant problem of Preference Optimisation for LLMs. Preference optimisation involves training a policy model to generate desirable outputs, often guided by a reward model that learns to distinguish between high-quality and low-quality outputs. This process, in many cases, entails iterative training of both models, generally in settings of online learning Guo et al. (2024a) or as a way to refine the reward model to account for the distribution shift of the policy throughout training Wang et al. (2024).
To this end, we hypothesise that the evolving dynamics of the policy and reward models during Preference Optimisation can provide valuable information for guiding the search direction of Zeroth-Order methods. We thus begin our investigation by visualising and analysing how popular preference optimisation methods, such as Proximal Policy Optimization (PPO) Schulman et al. (2017); Ouyang et al. (2022) and Direct Preference Optimization [^1] As standard DPO does not use any reward directly, we choose one of its variations Guo et al. (2024a) where we can integrate the reward model. (DPO) Rafailov et al. (2023), act on the policy and reward model, with particular attention on their relation. Whilst we do not find any statistical direct correlation, our visualisation and analysis suggest a significant influence of previous updates on upcoming ones. Further, we find that the reward model’s updates are consistently orthogonal to those of the policy.
@@
    Based on the observed dynamics, we develop ZOPrO, a Zeroth-Order Optimisation algorithm specifically tailored for preference optimisation in LLMs. ZOPrO incorporates a novel mechanism to guide the search direction by utilising the inter-dependencies between the policy and reward models.
@@
    We conduct experiments on multiple tasks and datasets to demonstrate the effectiveness of ZOPrO in fine-tuning LLMs for Preference Optimisation. Our results show that ZOPrO achieves comparable performance to first-order methods while significantly reducing memory usage and iteration time.
@@ 2.1 Preference Optimisation
Reinforcement Learning from Human Feedback (RLHF), or more generally, Preference Optimisation, aims to optimise Machine Learning models to align with complex human judgements and preferences Christiano et al. (2017). Typically, human feedback data is used to train a reward model, often conceptualised with the Bradley-Terry model Bradley and Terry (1952a), as an effective approach to converting human preferences into continuous reward signals (Bai et al., 2022). The reward model is then used to optimise a policy model using Reinforcement Learning algorithms, such as Proximal Policy Optimization (PPO) Schulman et al. (2017). RLHF has been applied in LLMs across various tasks from sentiment generation Ziegler et al. (2019) to summarisation Stiennon et al. (2020b) to instruction following Ouyang et al. (2022). In this framework, the policy and reward models influence each other in a shared interplay. Updates to the policy shift the text distribution that the reward model evaluates (Lu et al., 2024; Wang et al., 2024; Dou et al., 2024), while updates to the reward model alter the policy’s optimisation trajectory (Li et al., 2024).
Newer methods, such as Direct Preference Optimization (DPO) Rafailov et al. (2023) and variations Liu et al. (2024); Gheshlaghi Azar et al. (2024); Ji et al. (2024) avoid the use of Reinforcement Learning, instead relying on indirect reward. However, other works suggest that the use of a reward model or an external judge can still be beneficial Guo et al. (2024a). Whilst promising, these methods have not replaced RL-based approaches. Newer RL approaches include Group Relative Policy Optimization (GRPO) Shao et al. (2024), which introduces group-based relative rewards to accelerate convergence and obtain substantial improvements in math and general reasoning DeepSeek-AI (2025) or Ahmadian et al. (2024) which introduces REINFORCE Leave-One-Out (RLOO), a REINFORCE-based approach competitive to PPO with gains in memory utilisation and convergence speed.
Recent works have also explored operations in the weight space to enhance preference alignment. ExPO Zheng et al. (2025) proposes a linear extrapolation technique to improve model performance after partial or full preference optimisation. Related trends include Rewarded Soups Rame et al. (2023), which uses Linear Mode Connectivity (LMC) Frankle et al. (2020) to interpolate between models optimised on diverse reward signals. These methods rely on the assumption that preference-optimised models remain connected via simple geometric transformations, such as linear paths in parameter space. While compelling, this assumption is not general and may not (and as we will show, does not) hold in settings where policy and reward models evolve under distinct and potentially divergent optimisation pressures.
@@ 3 Visualising Preference Optimisation
We begin our investigation by analysing how the policy and reward models are optimised using traditional methods. We simulate multiple training runs across different methods, collecting model checkpoints at various stages of optimisation. At the end of training, we visualise and analyse the weights as updated by each method with particular attention to the interplay between the reward and policy models.
@@ 3.1 Experimental Setup
We select the two most used methods for preference optimisation: PPO Schulman et al. (2017) and DPO Rafailov et al. (2023); for the latter, we select its online version Guo et al. (2024a) which integrates a reward model—an essential component for our analysis. The models are trained using Anthropic’s HH-RLHF dataset Bai et al. (2022), chosen due to its widespread adoption in preference modelling tasks. After initial testing, we report convergence difficulties with PPO, where the model easily collapses into generating only eos tokens. Due to this, we instead analyse one of its variations, RLOO Ahmadian et al. (2024). RLOO is very similar to PPO in many aspects, but differs in how it treats actions and in the complexity of its objective. PPO treats each generated token as an individual action. It computes advantages with an extra value model [^2] Here note that even for PPO, we use the same model both as the value and reward for simplicity. using Generalized Advantage Estimation (GAE), and optimises a clipped policy ratio. RLOO simplifies this by treating the entire completion as a single action, avoiding the value model altogether. It uses a REINFORCE-style objective with a leave-one-out baseline, sampling $k$ ($=2$ in our setting) rollouts per query.
For all methods, we adopt an iterative training regime in which, after each iteration, the reward model is updated to better evaluate samples from the policy model, accounting for distribution shifts. We show pseudocode for our training regime in Algorithm 1. For the reward, we use the standard Bradley-Terry model Bradley and Terry (1952b).
@@ 3.3 Results
We show some 2D visualisations of different methods in Figure 3 and 3D visualisations in 2. We include interactive visualisations at https://alessiogalatolo.github.io/VisZOPrO/.
@@
These insights align with existing literature, where DPO has been shown to exhibit more stable training dynamics compared to RL-based approaches. On the other hand, the non-linear trajectory of optimisation goes in contrast with some literature on model merging Zheng et al. (2025), which attempts to obtain fully trained models by linear extrapolation of partial training weights.
@@ 3.4 Quantitative Analysis
Motivated by our observations on relative angles, we conduct a detailed analysis of orthogonality. Specifically, we examine model weights in their original dimensional space, assessing orthogonality both at the overall model level and layer-wise. Our empirical findings demonstrate consistent orthogonality between reward and policy updates across both methods. While both models originate from the same pretrained model (e.g., Llama), their updates quickly diverge due to distinct optimisation objectives. At iteration 0, this scenario closely mirrors that of Yu et al. (2020), who showed that conflicting gradient updates naturally arise in multi-objective contexts with shared initial parameters. As training progresses, however, the analogy weakens due to increasing parameter divergence. Nonetheless, our results extend the broader insight empirically into the context of two interacting but distinct processes—reward modelling, which employs discriminative supervised learning, and policy optimisation, which involves generative preference optimisation.
@@ 4.1 Formalisation
Our proposed ZOPrO method utilises this structure in conjunction with SPSA Spall (1992), as the base of the ZO algorithm, and RLOO Ahmadian et al. (2024), as the objective function.  [^4] Whilst we use RLOO as an exemplar objective, our method can be easily extended to other RL-based objectives without substantial changes.
@@ 4.1.1 Optimisation via SPSA
Note how our approach is also validated by previous attempts at surrogate or prior-guided strategies in ZO methods such as Cheng et al. (2021); Meier et al. (2019), which influence their sampling through the use of multiple ‘oracle gradient directions’. In our case, we ground our perturbation strategy in empirical evidence and reuse the updates from the previous iteration. In particular, prior works had to use ‘oracle’ directions as ZO updates are very noisy locally and cannot be used directly. Our method instead leverages the weight updates over a whole iteration, making their use possible.
@@ 4.1.2 Adaptive Perturbation Strategy
To compute perturbation directions $z$, we first utilise $\Delta\pi_{t}$ with a scaling factor $\alpha$, which starts at 1 and decays to 0 throughout each iteration. Next, we sample a random noise vector from a multivariate standard normal distribution, $\mathcal{N}(0,I_{d})$, and project it onto the orthogonal complement of $\Delta R_{t}$, yielding $u_{i}$. A preliminary perturbation vector $\hat{z}_{i}$ is given by:
@@ 4.2 Convergence Analysis
First, the deterministic component $\Delta\pi_{t}$ in our perturbation formulation vanishes as $\alpha$ decays to zero, ensuring that stochastic exploration remains dominant late in the iteration. This preserves the key stochastic approximation properties required for convergence.
@@ 5 Experiments
Among the challenges of effective optimisation of ZO methods is a careful balance of the hyperparameters $\eta$ and $\varepsilon$: when these values are too high, the model rapidly degenerates into generating incoherent outputs; if they are too low, the model fails to improve. The challenge is further exacerbated in the context of Preference Optimisation, where improvements are subtle compared to classification tasks (previous works), in which performance gains are more easily quantifiable. For this, we sweep over multiple values of $\varepsilon$ (the perturbation multiplier) and learning rates $\eta$ using a standard grid search approach where we check 9 combinations of these values. We select the most promising ones based on their reward improvement and keep them fixed across tasks.
For our experiments, we adopt a similar iterative training to that described in Section 3.1 and Algorithm 1. We initially only consider a single model, Llama 3.2 1B Llama Team (2024), the previously used task of building a conversational assistant and add two new tasks: summarisation, through the dataset ‘Summarize from Feedback’ Stiennon et al. (2020a), and machine translation through the dataset WMT20 Barrault et al. (2020).
@@ 5.1 Results
In Table 1 we show various metrics comparing our method with RLOO and PPO [^5] We do not compare with Online DPO due to the difference in objectives between the methods. over a complete iteration of training. Here, we report the average reward at the end of the iteration compared to the beginning, the time the method took to complete the iteration, the reward ‘gained’ over time (i.e., how fast in wall-clock time the method is improving) and the peak memory usage. We can see that our method whilst reporting the lowest reward ‘gains’ at the end of the iteration, is the fastest. This conforms to previous ZO methods where every step brings very little improvement but at much less cost in terms of time and memory.
@@
5.2 Hyperparameter Sensitivity
During our experiments, we found the values of $\varepsilon$ and $\eta$ to have a significant influence on training performance. Concretising in suboptimal training times and sometimes in policy degeneration. To further investigate this, we report the results of our hyperparameter search done with our method to obtain the results in Table 2. In particular, these results are relative to the ‘Conversation’ task that proved to be the most difficult during our testing.
Learning rate ($\eta$) Perturbation rate ($\varepsilon$) Reward $\uparrow$ 1e-3 1e-3 +0.06 1e-4 +0.2 1e-5 -1.9 (collapsed) 1e-4 1e-3 +0.1 1e-4 +0.14 1e-5 +0.03 1e-5 1e-3 +0.13 1e-4 +0.25 1e-5 +0.21
Table 2: Additional results on other model families. Gemma 2, contrary to the Llama and Qwen, does not officially support any language other than English.
Here, we note how even with a not particularly thorough hyperparameter search, most of the combinations yield good results, with more than half showing significant improvement. Based on this, we believe a good strategy for hyperparameter selection would be to select a small learning rate (e.g., 1e-5) and a perturbation that matches it. This should avoid model collapse at the cost of possibly slower improvement. Then, it is possible to speed up the performance gains by progressively increasing both hyperparameters, starting from the perturbation rate.
5.3 Additional Model Families
After initial positive results, we extend our experiments to two additional model families: Qwen 2.5 Qwen Team (2025) with its 1.5B model and Gemma 2 Gemma Team (2024) with its 2.4B model. These additional results, shown in Table 3, corroborate our methodology even further by showing consistent improvement across model families and tasks.
Model Task Reward $\uparrow$ Qwen 2.5 Conv. +0.28 Summ. +0.3 MT +0.15 Gemma 2 Conv. +0.13 Summ. +0.23 MT +0.25
Table 3: Additional results on other model families. Gemma 2, contrary to the Llama and Qwen, does not officially support any language other than English.